A 1 000-V battery, a 3 000-Ω resistor, and a 0.50-µF capacitor are connected in series with a switch. The capacitor is initially uncharged. What is the value of the current the moment after the switch is closed?
Answer #1

Initially, the uncharged capacitor acts like a short circuit immediately after the switch is closed i.e if the switch is closed at t = 0s, the capacitor acts like a short circuit at t=0+ s.

Value of the current the moment after the switch is closed = current in the circuit at t=0+ s = ( V/R) = (1000/ 3000) = (1/3) A (ans)